Let the set U = {1, 2, 3, 4, 5}, be the universe. Let Y = {(n+2)| n = 2, 3} 5 and X = {3, 4}. Write the following sets using set builder notation


1
Expert's answer
2022-04-26T10:07:37-0400

Set builder notation is a mathematical notation of describing a set by listing its elements or demonstrating its properties that its members must satisfy. Set "U" we can write in the form: "U=\{u: (u\in{\mathbb{N}})\land(1\leq u\leq 5)\}". "Y" can be rewritten in the following form: "Y=\{n:(n\in{\mathbb{N}})\land((n=4)\lor(n=5))\}". X can be rewritten in the following form: "X=\{x:(x\in{\mathbb{N}})\land((x=3)\lor(x=4))\}". In case we assume that "U" is the universe and we consider only subsets of "U", we get: "Y=\{n\in U: (n=4)\lor(n=5)\}" and "X=\{x\in U: (x=3)\lor(x=4)\}".